Could someone please check my work and show me how to finish this problem? solve by substitution
7m+n=42, m-7n=56
7m+n=42
7m=42-n
m=(42n)/7
m-7n=56
42-n-49=392

Answer by rfer(16322)

m-7n=56 m=7n+56
7m+n=42
7(7n+56)+n=42
49n+392+n=42
50n=-350
n=-7
---------------
7m-7=42
7m=49
m=7
----------------
proof
m-7n=56
7-7*-7=56
7+49=56